OpenStudy (anonymous):

solve for u. -6/u-5=3/2u-10+2 of there is more than one solution, separate them with commas. If there is no solution, put no solution

OpenStudy (anonymous):

\[-\frac{ 6 }{ u-5 }=\frac{ 3 }{ 2u-10}+2\]

OpenStudy (nikato):

can u factor the 2u-10

OpenStudy (anonymous):

2(u-5)

OpenStudy (nikato):

ok. now multiply the first fraction by 2/2. and multiply the 2 by 2(u-5)/2(u-5)

OpenStudy (anonymous):

wait so u want me to multiply -6/u-5?

OpenStudy (nikato):

 so they all have common denominator